a ball is dropped 3 meters. after the ball bounces, it raises to a height of 2 meters. the mass of the ball is 0.5 kg. What is the speed of the ball right before the bounce?
Answer #1

When the ball is dropped from a height h, initial potential energy is mgh You would expect it to accelerate to a velocity right before the bounce is :-

v=\sqrt{2gh}

Given is :-

h=3m,g=9.8m/s^2

by putting all the given values in above equation, we get

v=\sqrt{2 \times 9.8 \times 3} =7.66 m/s

Thus the speed of the ball right before the bounce is 7.66m/s